OpenStudy (anonymous):

Which binomial is a factor of 4x2-13+3?

OpenStudy (anonymous):

the last number is mutip of 3*1=3 but the middle is 13 which only (4*3) 12+1 (4*1= 4+12=16 ) or 10+2 (no possible with 3), in the end we need ot use the quadratic formula x=-b+-sqrt(b^2-4ac)/2a

OpenStudy (anonymous):